(1 point) A particle moves with its position given by x = cos(8t) and y = sin(t), where positions are given in feet from the I've tried like 1000000 times but I cannot find the speed of the particle... plz help me plz
Answer #1

of particle and somet y = sin(t) dy = CO3(+) Positions de = Cop (8t) 1 dx = - Bin (8t)8 - 8 Sim (8+) relocidy = plats o de do